# Smoothed Analysis of Online Metric Matching with a Single Sample: Beyond Metric Distortion

### v1

@@ 2.2 Algorithm with Sample Access to Request Distributions
[YY25] show that given an arbitrary algorithm $\mathcal{A}$ and one sample from each request distribution, we can construct another algorithm whose cost can be decomposed into the offline optimum and the cost of $\mathcal{A}$ when the server set is also drawn from $\mathbb{D}$. [^2] They only prove this statement for i.i.d. requests, and the generalization to the case with non-identical request distributions follows verbatim.
Lemma 2.4 (Theorem 2 in [YY25]).
@@ 3 Random-Subtree Algorithm
The Random-Subtree (RS) algorithm proposed by [GL12] is a natural starting point for online matching on HST metrics. In the adversarial setting, it achieves an $O(\log\Delta)$ competitive ratio for $\Delta$-ary $\alpha$-HST metrics, provided $\alpha=\Omega(\log\Delta)$. Here we revisit RS in the stochastic setting where servers and requests are independently drawn from a distribution $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$. Later on, we will translate the guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.4. Rather than its competitive ratio, we upper bound the cost of the $\mathrm{RS}$ algorithm for arbitrary $\alpha\geq 2$.
@@
For any $\Delta$-ary $\alpha$-HST metric with height $h$ and $\alpha\geq 2$, and for any $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$,
@@
The proof of Theorem 3.1 relies on the following worst-case guarantee for the $\mathrm{RS}$ algorithm, whose cost is upper bounded in terms of the number of excess requests in each subtree.
@@
For any $\Delta$-ary $\alpha$-HST metric with height $h$ and $\alpha\geq 2$, and for all $S$ and $R$,
@@
$\mathbb{E}[(\hat{r}(v)-\hat{s}(v))^{+}]\leq\mathbb{E}[|\hat{r}(v)-\mathbb{E}\hat{r}(v)|]+\mathbb{E}[|\hat{s}(v)-\mathbb{E}\hat{s}(v)|]=2\mathbb{E}[|\hat{r}(v)-\mathbb{E}\hat{r}(v)|].$ (3)
Moreover,
$\mathbb{E}[|\hat{r}(v)-\mathbb{E}\hat{r}(v)|]\leq\mathrm{std}(\hat{r}(v))=\mathrm{std}(\mathrm{PB}(\mu_{\mathbb{D}_{1}}(v),\ldots,\mu_{\mathbb{D}_{n}}(v)))\leq\sqrt{\sum_{i=1}^{n}\mu_{\mathbb{D}_{i}}(v)}=\sqrt{\mu_{\mathbb{D}}(v)},$ (4)
where the first inequality holds by Jensen’s inequality. As a result, for each $j\in\{0,1,\ldots,h-1\}$,
$\sum_{v\in V_{j}}\mathbb{E}[(\hat{r}(v)-\hat{s}(v))^{+}]\leq 2\sum_{v\in V_{j}}\mathbb{E}[|\hat{r}(v)-\mathbb{E}\hat{r}(v)|]\leq 2\sum_{v\in V_{j}}\sqrt{\mu_{\mathbb{D}}(v)}\leq 2\sqrt{n|V_{j}|}=2\sqrt{n\Delta^{h-j}},$ (5)
where the first inequality follows from (3), the second inequality follows from (4), and the third inequality holds by the concavity of $f(t)=\sqrt{t}$ and the fact that $\sum_{v\in V_{j}}\mu_{\mathbb{D}}(v)=n$.
@@
The proof of Lemma 3.2 largely follows the strategy of [GL12, Theorem 4.1], with two key differences. First, their analysis assumes $\alpha=\Omega(H_{\Delta})$, ensuring the HST is sufficiently well-separated so that costs incurred at lower levels can be absorbed into those at higher levels. In contrast, we do not rely on this separation, and hence the costs incurred at lower levels must be handled explicitly. Second, our proof is relatively simpler: since we only bound the absolute cost of the algorithm, we avoid the more involved step of characterizing the offline optimum, which is required in [GL12].
@@
Lemma 3.3 (Lemma 4.3 in [GL12]).
@@
We recall the following useful decomposition of $\mathrm{cost}(\mathrm{RS};S,R\cup R^{\prime})$ from [GL12].
Lemma 3.4 (Fact 4.8 in [GL12]).
@@ 4 BBGN Algorithm
In this section, we turn from the $\mathrm{RS}$ algorithm to the algorithm proposed by [BBGN14], henceforth referred to as the BBGN algorithm, which outperforms the $\mathrm{RS}$ algorithm in certain parameter regimes. The BBGN algorithm is $O(\log n)$-competitive for $\alpha$-HST metrics with any constant $\alpha>1$ when both $S$ and $R$ are adversarial. Our interest, however, lies in analyzing its cost when both $S$ and $R$ are drawn from $\mathbb{D}$. As before, we will later translate these guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.4.
For every $j\in[h]$, where $h$ is the height of the HST, define $V_{j}^{\prime}:=\{v\in V_{j}\mid\hat{s}(v)\geq 2\}$ to be the set of nodes with height $j$ whose subtrees have at least two servers. We now state the main theorem of this section.
@@
For any $\Delta$-ary $\alpha$-HST metric with height $h$ and constant $\alpha\geq 2$ such that $\Delta^{h-1}\leq n/{\color[rgb]{1,0,0}2}$, and for any $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$,
@@
The rest of this section is devoted to proving Theorem 4.1. For each leaf node $v$ of the HST and $\ell\in\{0,1,\ldots,h\}$, let $p(v,\ell)$ be the ancestor of $v$ with height $\ell$. The $\mathrm{BBGN}$ algorithm was originally developed in a restricted reassignment online model, which permits limited reassignment of previously arrived requests, and then adapted to the standard online model. We recall here a worst-case upper bound for the cost of the $\mathrm{BBGN}$ algorithm, obtained by combining [BBGN14, Lemmas 4.2, 4.3, and 4.4], and we refer to [BBGN14] for a more formal description of the algorithm.
Lemma 4.2 ([BBGN14]).
Consider an $\alpha$-HST metric with $\alpha\geq 2$ being a constant. For all $S$ and $R$, in addition to the resulting (random) matching $M$ between $S$ and $R$, the $\mathrm{BBGN}$ algorithm also produces a (random) min-cost perfect matching $M_{\mathrm{OPT}}$ between $S$ and $R$ such that
@@
It is well known that min-cost perfect matchings in HST metrics admit a useful characterization (see, e.g., [BBGN14, Lemma 4.1]): every min-cost perfect matching between $S$ and $R$ can be obtained by a greedy algorithm that repeatedly matches the closest request-server pair and then recurses on the remaining instance. Equivalently, the number of matched pairs whose two endpoints lie in different subtrees of a given internal node can be expressed as follows.
@@
For any HST metric, in any min-cost perfect matching between $S$ and $R$, the number of matches whose two endpoints lie in different subtrees of an internal node $v$ equals
@@
Combining Lemma 4.2 and 4.3, we can then upper bound the expected cost of the $\mathrm{BBGN}$ algorithm.
@@
For any $\alpha$-HST metric with $\alpha\geq 2$ being a constant, and for all $S$ and $R$,
$\mathrm{cost}(\mathrm{BBGN};S,R)\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sum_{i=1}^{\Delta}(\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v)))^{+}\right).$
@@
$=O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\cdot q(v)\right)$
$\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sum_{i=1}^{\Delta}(\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v)))^{+}\right),$ (9)
@@
$\mathrm{cost}(\mathrm{BBGN};\mathbb{D},\mathbb{D})$ $\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sum_{i=1}^{\Delta}\mathbb{E}[(\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v)))^{+}]\right)$
$\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sum_{i=1}^{\Delta}\sqrt{\mu_{\mathbb{D}}(c_{i}(v))}\right),$ (10)
where the second inequality follows from (3) and (4).
Since $\sum_{i=1}^{\Delta}\mu_{\mathbb{D}}(c_{i}(v))=\mu_{\mathbb{D}}(v)$ for every internal node $v$, by the concavity of $f(t)=\sqrt{t}$,
$\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sum_{i=1}^{\Delta}\sqrt{\mu_{\mathbb{D}}(c_{i}(v))}\leq\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\cdot\Delta\sqrt{\frac{\mu_{\mathbb{D}}(v)}{\Delta}}=\sqrt{\Delta}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sqrt{\mu_{\mathbb{D}}(v)}$ (11)
for every $j\in[h]$. Since $\sum_{v\in V_{j}^{\prime}}\mu_{\mathbb{D}}(v)\leq\sum_{v\in V_{j}}\mu_{\mathbb{D}}(v)=n$, and $\hat{s}(v)\geq 2$ for every $v\in V_{j}^{\prime}$, we apply again the concavity of $f(t)=\sqrt{t}$ to obtain
$\sqrt{\Delta}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sqrt{\mu_{\mathbb{D}}(v)}\leq\sqrt{\Delta}\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sqrt{\frac{n\log^{2}(\hat{s}(v))}{\sum_{u\in V_{j}^{\prime}}\log^{2}(\hat{s}(u))}}=\sqrt{n\Delta}\cdot\sqrt{\sum_{v\in V_{j}^{\prime}}\log^{2}(\hat{s}(v))}.$ (12)
Moreover, since $\sum_{v\in V_{j}^{\prime}}\hat{s}(v)\leq\sum_{v\in V_{j}}\hat{s}(v)=n$ and $\hat{s}(v)\geq 2$ for every $v\in V_{j}^{\prime}$, by the concavity of $f(t)=\log^{2}t$ over $e,+\infty)$,
$\sqrt{\sum_{v\in V_{j}^{\prime}}\log^{2}(\hat{s}(v))}$ $\leq\sqrt{\sum_{v\in V_{j}^{\prime}}\log^{2}(2\hat{s}(v))}\leq\sqrt{\sum_{v\in V_{j}^{\prime}}\log^{2}\frac{2n}{|V_{j}^{\prime}|}}$
$=\sqrt{|V_{j}^{\prime}|}\cdot\log\frac{2n}{|V_{j}^{\prime}|}\leq\sqrt{|V_{j}|}\cdot\log\frac{2n}{|V_{j}|}=\sqrt{\Delta^{h-j}}\cdot\log\frac{2n}{\Delta^{h-j}},$ (13)
where the last inequality holds since $f(t)=t\log^{2}(2n/t)$ is monotonically increasing over $(0,n/2]$, and $|V_{j}|=\Delta^{h-j}\leq\Delta^{h-1}\leq n/2$. Combining (11), (12), and (4), it follows that
$\sum_{v\in V_{j}^{\prime}}\log(\hat{s}(v))\sum_{i=1}^{\Delta}\sqrt{\mu_{\mathbb{D}}(c_{i}(v))}\leq\sqrt{n\Delta^{h-j+1}}\cdot\log\frac{2n}{\Delta^{h-j}}$ (14)
for every $j\in[h]$.
Finally, combining (10) and (14) concludes the proof. ∎

### v2

@@ 2.2 Algorithm with Sample Access to Request Distributions
[yang2025online] show that given an arbitrary algorithm $\mathcal{A}$ and one sample from each request distribution, we can construct another algorithm whose cost can be decomposed into the offline optimum and the cost of $\mathcal{A}$ when the server set is also drawn from $\mathbb{D}$. [^2] They only prove this statement for i.i.d. requests, but the generalization to the case with non-identical request distributions follows verbatim.
Lemma 2.5 (Theorem 2 in [yang2025online]).
@@ 3 Random-Subtree Algorithm
The Random-Subtree (RS) algorithm proposed by [DBLP:conf/icalp/GuptaL12] is a natural starting point for online matching on HST metrics. In the adversarial setting, it achieves an $O(\log\Delta)$ competitive ratio for $\Delta$-ary $\alpha$-HST metrics, provided $\alpha=\Omega(\log\Delta)$. Here we revisit the RS algorithm in the stochastic setting where servers and requests are independently drawn from a distribution $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$. Later on, we will translate the guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.5. Rather than its competitive ratio, we upper bound the cost of the $\mathrm{RS}$ algorithm for arbitrary $\alpha\geq 2$.
@@
For any $\Delta$-ary $\alpha$-HST with height $h$ and $\alpha\geq 2$, and for any $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$,
@@
The proof of Theorem 3.1 relies on the following bound for the $\mathrm{RS}$ algorithm, whose cost is upper bounded in terms of the number of excess requests in each subtree.
@@
For any $\Delta$-ary $\alpha$-HST with height $h$ and $\alpha\geq 2$, and for all $S$ and $R$,
@@
$\mathbb{E}[(\hat{r}(v)-\hat{s}(v))^{+}]\leq\mathbb{E}[|\hat{r}(v)-\mathbb{E}\hat{r}(v)|]+\mathbb{E}[|\hat{s}(v)-\mathbb{E}\hat{s}(v)|]=2\mathbb{E}[|\hat{r}(v)-\mathbb{E}\hat{r}(v)|]\leq 2\cdot\mathrm{std}(\hat{r}(v)),$ (3)
where the last inequality holds by Jensen’s inequality. Moreover,
$\mathrm{std}(\hat{r}(v))=\mathrm{std}(\mathrm{PB}(\mu_{\mathbb{D}_{1}}(v),\ldots,\mu_{\mathbb{D}_{n}}(v)))\leq\sqrt{\sum_{i=1}^{n}\mu_{\mathbb{D}_{i}}(v)}=\sqrt{\mu_{\mathbb{D}}(v)}.$ (4)
As a result, for each $j\in\{0,1,\ldots,h-1\}$,
$\sum_{v\in V_{j}}\mathbb{E}[(\hat{r}(v)-\hat{s}(v))^{+}]\leq 2\sum_{v\in V_{j}}\sqrt{\mu_{\mathbb{D}}(v)}\leq 2\sqrt{n|V_{j}|}=2\sqrt{n\Delta^{h-j}},$ (5)
where the first inequality follows from (3) and (4), and the second inequality holds by the concavity of $f(t)=\sqrt{t}$ and the fact that $\sum_{v\in V_{j}}\mu_{\mathbb{D}}(v)=n$.
@@
The proof of Lemma 3.2 largely follows the proof strategy of [DBLP:conf/icalp/GuptaL12, Theorem 4.1], with two key differences. First, their analysis assumes $\alpha=\Omega(H_{\Delta})$, ensuring the HST is sufficiently well-separated so that the costs incurred at lower levels are dominated by those at higher levels. In contrast, we do not rely on this separation, and hence the costs incurred at lower levels must be handled explicitly. Second, our proof is relatively simpler: since we only bound the absolute cost of the algorithm, we avoid the more involved step of characterizing the offline optimum, which is required in [DBLP:conf/icalp/GuptaL12].
@@
Lemma 3.3 (Lemma 4.3 in [DBLP:conf/icalp/GuptaL12]).
@@
We recall the following useful decomposition of $\mathrm{cost}(\mathrm{RS};S,R\cup R^{\prime})$ from [DBLP:conf/icalp/GuptaL12].
Lemma 3.4 (Fact 4.8 in [DBLP:conf/icalp/GuptaL12]).
@@ 4 BBGN Algorithm
In this section, we turn from the $\mathrm{RS}$ algorithm to the algorithm proposed by [DBLP:journals/algorithmica/BansalBGN14], henceforth referred to as the BBGN algorithm, which outperforms the $\mathrm{RS}$ algorithm in certain parameter regimes. The BBGN algorithm is $O(\log n)$-competitive for $\alpha$-HST metrics with any constant $\alpha>1$ when both $S$ and $R$ are adversarial. Our interest, however, lies in analyzing its cost when both $S$ and $R$ are drawn from $\mathbb{D}$. As before, we will later translate these guarantees to the setting where servers are adversarial and requests are drawn from $\mathbb{D}$ by applying Lemma 2.5.
We now state the main theorem of this section.
@@
For any $\Delta$-ary $\alpha$-HST with height $h$ and constant $\alpha\geq 2$ such that $\Delta^{h-1}\leq n/2$, and for any $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$,
@@
The rest of this section is devoted to proving Theorem 4.1. For each leaf node $v$ of the HST and $\ell\in\{0,1,\ldots,h\}$, let $p(v,\ell)$ be the ancestor of $v$ with height $\ell$. The $\mathrm{BBGN}$ algorithm was originally developed in a restricted reassignment online model, which permits limited reassignment of previously arrived requests, and then adapted to the standard online model. We recall here a bound for the cost of the $\mathrm{BBGN}$ algorithm that applies to all $S$ and $R$, obtained by combining [DBLP:journals/algorithmica/BansalBGN14, Lemmas 4.2, 4.3, and 4.4], and we refer to [DBLP:journals/algorithmica/BansalBGN14] for a more formal description of the algorithm.
Lemma 4.2 ([DBLP:journals/algorithmica/BansalBGN14]).
Consider an $\alpha$-HST with $\alpha\geq 2$ being a constant. For all $S$ and $R$, in addition to the resulting (random) matching $M$ between $S$ and $R$, the $\mathrm{BBGN}$ algorithm also produces a (random) min-cost perfect matching $M_{\mathrm{OPT}}$ between $S$ and $R$ such that
@@
It is well known that min-cost perfect matchings in HSTs admit a useful characterization (see, e.g., [DBLP:journals/algorithmica/BansalBGN14, Lemma 4.1]): every min-cost perfect matching between $S$ and $R$ can be obtained by a greedy algorithm that repeatedly matches the closest request-server pair and then recurses on the remaining instance. Equivalently, the number of matched pairs whose two endpoints lie in different subtrees of a given internal node can be expressed as follows.
@@
For any HST, in any min-cost perfect matching between $S$ and $R$, the number of matches whose two endpoints lie in different subtrees of an internal node $v$ equals
@@
Combining Lemma 4.2 and 4.3, we can then upper bound the expected cost of the $\mathrm{BBGN}$ algorithm as in the following lemma.
@@
For any $\alpha$-HST with $\alpha\geq 2$ being a constant, and for all $S$ and $R$,
$\mathrm{cost}(\mathrm{BBGN};S,R)\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}}\log(\hat{s}(v)+e)\sum_{i=1}^{\Delta}|\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v))|\right).$
@@
$\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}}\log(\hat{s}(v)+e)\cdot q(v)\right)$
$\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}}\log(\hat{s}(v)+e)\sum_{i=1}^{\Delta}|\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v))|\right),$ (9)
@@
$\mathrm{cost}(\mathrm{BBGN};\mathbb{D},\mathbb{D})$
$\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}}\sum_{i=1}^{\Delta}\mathbb{E}[\log(\hat{s}(v)+e)\cdot|\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v))|]\right)$
$\leq O\left(\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}}\sqrt{\mathbb{E}[\log^{2}(\hat{s}(v)+e)]}\sum_{i=1}^{\Delta}\sqrt{\mathbb{E}[(\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v)))^{2}]}\right),$ (10)
where the second inequality holds by the Cauchy-Schwarz inequality.
We first bound the last summation in (10). For all internal node $v$ and $i\in[\Delta]$,
$\sqrt{\mathbb{E}[(\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v)))^{2}]}$ $\leq\sqrt{2\mathbb{E}[(\hat{r}(c_{i}(v))-\mu_{\mathbb{D}}(c_{i}(v)))^{2}+(\hat{s}(c_{i}(v))-\mu_{\mathbb{D}}(c_{i}(v)))^{2}]}$
$=2\cdot\mathrm{std}(\hat{r}(c_{i}(v)))\leq 2\sqrt{\mu_{\mathbb{D}}(c_{i}(v))},$
where the first inequality follows since $(x+y)^{2}\leq 2(x^{2}+y^{2})$, the equality holds since $\hat{r}(c_{i}(v))$ and $\hat{s}(c_{i}(v))$ are identically distributed with $\mathbb{E}[\hat{r}(c_{i}(v))]=\mathbb{E}[\hat{s}(c_{i}(v))]=\mu_{\mathbb{D}}(c_{i}(v))$, and the last inequality holds by (4). It follows that
$\sum_{i=1}^{\Delta}\sqrt{\mathbb{E}[(\hat{r}(c_{i}(v))-\hat{s}(c_{i}(v)))^{2}]}\leq 2\sum_{i=1}^{\Delta}\sqrt{\mu_{\mathbb{D}}(c_{i}(v))}\leq 2\sqrt{\Delta\cdot\mu_{\mathbb{D}}(v)},$ (11)
where the last inequality holds by the concavity of $f(t)=\sqrt{t}$ and the fact that $\sum_{i=1}^{\Delta}\mu_{\mathbb{D}}(c_{i}(v))=\mu_{\mathbb{D}}(v)$.
Next, we bound $\mathbb{E}[\log^{2}(\hat{s}(v)+e)]$ for every internal node $v$. Since $f(t)=\log^{2}(t+e)$ is concave over $0,+\infty)$, by Jensen’s inequality,
$\mathbb{E}[\log^{2}(\hat{s}(v)+e)]\leq\log^{2}(\mathbb{E}[\hat{s}(v)]+e)=\log^{2}(\mu_{\mathbb{D}}(v)+e).$ (12)
Finally, combining (10), (11), and (12),
$\mathrm{cost}(\mathrm{BBGN};\mathbb{D},\mathbb{D})$ $\leq O\left(\sqrt{\Delta}\sum_{j=1}^{h}\alpha^{j-h}\sum_{v\in V_{j}}\log(\mu_{\mathbb{D}}(v)+e)\sqrt{\mu_{\mathbb{D}}(v)}\right)$
$\leq O\left(\sqrt{\Delta}\sum_{j=1}^{h}\alpha^{j-h}\cdot|V_{j}|\cdot\sqrt{\frac{n}{|V_{j}|}}\log\left(\frac{n}{|V_{j}|}+e\right)\right)$
$\leq O\left(\sqrt{n\Delta}\sum_{j=1}^{h}\alpha^{j-h}\sqrt{\Delta^{h-j}}\log\frac{n}{\Delta^{h-j}}\right),$
where the second inequality holds by the concavity of $f(t)=\sqrt{t}\log(t+e)$ over $0,+\infty)$ and the fact that $\sum_{v\in V_{j}}\mu_{\mathbb{D}}(v)=n$, and the third inequality holds since $|V_{j}|=\Delta^{h-j}\leq n/2$ for every $j\in[h]$. This concludes the proof. ∎